# AlphaForge: A Framework to Mine and Dynamically Combine Formulaic Alpha Factors

### v1

@@ Abstract.
The variability and low signal-to-noise ratio in financial data, combined with the necessity for interpretability, make the alpha factor mining workflow a crucial component of quantitative investment. Transitioning from early manual extraction to genetic programming, the most advanced approach in this domain currently employs reinforcement learning to mine a set of combination factors with fixed weights. However, the performance of resultant alpha factors exhibits inconsistency, and the inflexibility of fixed factor weights proves insufficient in adapting to the dynamic nature of financial markets. To address this issue, this paper proposes a two-stage formulaic alpha generating framework AlphaForge, for alpha factor mining and factor combination. This framework employs a generative-predictive neural network to generate factors, leveraging the robust spatial exploration capabilities inherent in deep learning while concurrently preserving diversity. The combination model within the framework incorporates the temporal performance of factors for selection and dynamically adjusts the weights assigned to each component alpha factor. Experiments conducted on real-world datasets demonstrate that our proposed model outperforms contemporary benchmarks in formulaic alpha factor mining. Furthermore, our model exhibits a notable enhancement in portfolio returns within the realm of quantitative investment.
Computational Finance, Quantitative investment, Stock Trend Forecasting, Alpha Mining
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: ; 2024; Beijing
1. Introduction
A central challenge in the field of quantitative investment is stock trend forecasting. This difficulty arises from the inherent characteristics of stock data, featuring a low signal-to-noise ratio and considerable noise (Qian et al., 2007). Professionals and researchers commonly employ a strategy involving the extraction of Alpha factors from raw historical data to predict future returns.
@@ 1. Introduction
The contemporary approaches to Alpha factor mining can be broadly categorized into two main methodologies: machine learning methods and formulaic Alpha methods. Deep learning models such as Long Short-Term Memory networks LSTM(Hochreiter and Schmidhuber, 1997), Historical Inference Sequence Transformers (HIST)(Xu et al., 2021a), VectorVAE(Duan et al., 2022), Recursive Ensemble of Small Trees (REST)(Xu et al., 2021b), among others, are commonly employed to generate more complex Alphas, which often lack interpretability, operating essentially as black-box functions with inputs and outputs. On the other hand, formulaic Alpha mining aims to identify simple formulas that can replace Alpha factors. Early approaches centered on the manual extraction of factors with economic significance, such as the Fama three-factor model(Fama and French, 1992), and Fama five-factor model. However, as these factors became widely known and used, their ability to predict outcomes started to decrease(Kakushadze, 2016). Over time, to overcome the limitations of traditional factors, researchers began exploring alternative methods, such as genetic programming, to automatically generate more effective factors. (Kakushadze, 2016) conducts an analysis of 101 formulaic Alpha factors within the US market. Employing a genetic programming algorithm, the study systematically explores individual, unrelated factors through genetic variation of formula trees. The current state-of-the-art method involves using reinforcement learning algorithms to simultaneously identify a combination of Alpha factors along with their associated weights(Yu et al., 2023), with the objective of optimizing the discovery of the most robust composite factors.
@@
In investment practice, investment managers often collect a large batch of Alpha factors into a factor library. These alpha factors are combined through a combination model to form ”Mega-Alpha”, a final signal used for trading. With each new data arrival, factor values are recalculated, and Mega-Alpha is computed for trading decisions. Due to the financial market’s emphasis on interpretability, the models for combining factors typically adopt linear structures. Cutting-edge Reinforcement Learning (RL) methods(Yu et al., 2023) integrate the combination model and mining process into a unified framework, determining both factor discovery and their combination into Mega-Alpha. However, despite extensive factor exploration, only a fixed subset is typically utilized in actual investment. Additionally, the cyclic nature of each Alpha factor’s stock selection capability, and the potential for reversal over time, must be considered. Fixed factors and weights sometimes render parts of the Mega-Alpha ineffective or even reverse their effects.
@@
The factor mining model, inspired by (Linder et al., 2020), utilizes a surrogate model to learn the distribution of alpha factor scores. The generative model is trained to maximize the output of the surrogate model, facilitating the generation of high-scoring factors. Gradient-based methods ensure the generation of desirable factors even in the extremely sparse score space, allowing adjustments to the scoring function based on previously mined factors to ensure low correlation.
@@
  3. (3)
    We conducted a comprehensive set of experiments to validate the effectiveness of our proposed methodology. Subsequent additional experiments provided evidence that factor timing can result in profitable outcomes.
2. Preliminary
2.1. Alpha Factor Definition
@@
2.2. Alpha Factor Metrics
@@
(1) $IC(f,X,Y)=\mathbb{E}_{t}\left[\rho\left(v_{t},y_{t}\right)\right]=\frac{1}{T}\sum_{t=1}^{T}\rho\left(v_{t},y_{t}\right)$
where $v_{t},y_{t}\in\mathbb{R}^{n}$. The correlation $\rho\left(v_{t},y_{t}\right)$ for each cross-section depicts the relationship between the factor value and the subsequent period’s return. The IC describes the overall stock-picking ability of the factor, with higher values indicating superior stock-picking performance. However, a factor with a high IC may exhibit unstable stock-picking performance, particularly if it experiences substantial daily fluctuations in its IC. In such cases, despite a high IC, the factor cannot be considered an excellent alpha factor, as it introduces volatility in returns. Consequently, the ICIR (Information Coefficient Information Ratio) metric becomes imperative:
(2) $\operatorname{ICIR}(f,X,Y)=\frac{\mathbb{E}_{t}\left[\rho\left(v_{t},y_{t}\right)\right]}{\operatorname{Std}\left(\rho\left(v_{t},y_{t}\right)\right)}$
The ICIR metric, which assesses the stock-picking ability of the factor while incorporating considerations of its stability, is deemed more favorable with higher values. Since the denominator of ICIR, $\operatorname{Std}\left(\rho\left(v_{t},y_{t}\right)\right)\geq 0$ , the sign of ICIR coincides with that of IC.
Pearson correlation is easily influenced by outliers, potentially leading to deviations from the true correlation in the final results. Consequently, the inclusion of RankIC and RankICIR is necessary to complement the measurement indicators:
(3) $\operatorname{RankIC}(f,X,Y)=\mathbb{E}_{t}\left[\tilde{\rho}\left(v_{t},y_{t}\right)\right]=\frac{1}{T}\sum_{t=1}^{T}\tilde{\rho}\left(v_{t},y_{t}\right)$
(4) $\operatorname{RankICIR}(f,X,Y)=\frac{\mathbb{E}_{t}\left[\tilde{\rho}\left(v_{t},y_{t}\right)\right]}{\operatorname{Std}t\left[\tilde{\rho}\left(v_{t},y_{t}\right)\right]}$
where $\tilde{\rho}\left(v_{t},y_{t}\right)$ is the Spearman’s rank correlation coefficient. Similar to IC and ICIR, higher values of RankIC and RankICIR are also better, and the sign of RankICIR aligns with that of RankIC. It is noteworthy that IC, ICIR, Rank_IC, and RankICIR all have the following characteristics:
(5) $\mathcal{F}(f,X,Y)=-\mathcal{F}(-f,X,Y)\quad$
This implies that when a factor’s IC (Information Coefficient) is negative, it is possible to reverse the sign of the factor value to rectify the original factor into one with a positive IC. Consequently, during the factor mining process, the absolute value of the IC can be adopted as the mining objective.
2.3. Formulaic Alpha
@@
The representation of formulas traditionally involves expression trees, where operands are leaf nodes, and operators are non-leaf nodes. Initial attempts at formula generation employed genetic programming algorithms on these trees. In alignment with contemporary deep learning methods, our approach follows the methodology introduced in (Yu et al., 2023). This involves representing formulas using Reverse Polish Notation (RPN), acquired through post-order traversal of the formula tree. This facilitates the representation of formulas in a one-hot matrix format compatible with deep learning networks.
3. Methodology
@@
3.1. Factor Mining Model
Figure 1. The Demonstration of Factor Generating Model
@@ 3.1. Factor Mining Model
$\mathcal{L}_{P}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left(P\left(\boldsymbol{x}_{i}\right)-\text{ fitness }\left(\boldsymbol{x}_{\boldsymbol{i}}\right)\right)^{2}}.$
@@
$\mathcal{L}_{\text{Fitness }}=-P(M(G(z)))$
@@
However, focusing solely on optimizing for high fitness may lead to premature convergence of the network $G$ to a local optimum. Therefore, it is necessary to introduce a diversity loss to force the generator $G$ to produce a diverse array of alpha factor formulas. To achieve this, we introduce a loss by generating two sets of factors based on two samplings of $z_{1}$ and $z_{2}$, subsequently penalizing the correlation between these two sets of factors: {dmath} L_Diversity =λ_onehot * Similarity _onehot (M(G(z_1)), M(G(z_2)))+λ_hidden * Similarity_hidden ((M(G(z_1)), M(G(z_2))) The final loss function for training the generator $\mathrm{G}$ is: {dmath} L_G=L_Fitness +L_Diversity =-P(x_1)+λ_onehot * Similarity _onehot (f(z_1), f(z_2))+λ_hidden * Similarity _hidden (f(z_1), f(z_2))
@@
Our framework is designed to identify a set of highly effective strong factors in the factor mining stage for the factor timing combination model. This involves an investigation, including a comparison between strong and weak factor pools. A strict criterion is maintained to ensure that factors included in the library meet specific requirements. Incorporating domain knowledge, our criteria for factor entry comprise three fundamental aspects:Information Coefficient (IC), IC Information Ratio (ICIR), and the correlation of returns with existing factors in the library. IC serves as an indicator of the factor’s stock-picking ability, ICIR signifies the stability of this ability, and return correlation ensures that the factor avoids overlapping with the stock-picking ability of existing factors. Due to the implementation of a generative-predictive architecture coupled with gradient-based algorithms, the generator is capable of capturing the essential characteristics of factors in a ”directional” manner, even under conditions of significant sparsity within the fitness function. The fitness score function is defined as follows: {dmath} π(x, Z, X, Y)= { Abs (IC(f, X, Y)),f is valid and returncorr(f, Z, X, Y)¡CORR0, else Where $f=\operatorname{parse}(x)$ represents the operational formula parsed from the onehot matrix representation $x$, and $\mathcal{Z}$ denotes the existing factor set which is called factor zoo. The absolute value of IC is taken because a negative IC factor can be transformed into a positive IC factor by reversing the value of the factor as shown in Equation 5. As long as the absolute value of the factor IC is sufficiently high, it is considered a good factor. The $\operatorname{returncorr}$ function computes the maximum absolute value of the correlation between $f$ and the return rate of each existing factor in $\mathcal{Z}$. Here, CORR’ is a manually set parameter. When $|Z|=0$, $\pi(f,\mathcal{Z},X,Y)$ returns the absolute value of the factor’s IC. It is important to note that as the number of factors in the factor zoo $Z$ increases, the potential number of factors meeting the condition $\operatorname{returncorr}(x,Z,X,Y)<\operatorname{CORR}$ ’ will decrease sharply. Consequently, fitness scores for most factors will be set to 0, resulting in a sparse exploration space and an increase in the difficulty of factor mining. Algorithm 1 provides a detailed description of our alpha mining model.
Algorithm 1 Factor Mining Pipeline
1:Input: stock data including data and target $X=\{X_{t}\},Y=\{y_{t}\}$
2:Output: A group of respectively low correlation strong factors which is called factor zoo: $\mathcal{Z}=\{f_{1},\ldots,f_{k}\}$
3:Initialize the factor zoo $\mathcal{Z}=\emptyset$
4:Sample a group of randomized factors onehot matrices =$\mathcal{R}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{r}\}$
5:while $|\mathcal{Z}|<\text{TargetFactorNum}$ do
6:     $\mathcal{R}_{\text{fitness}}=\{\pi(\mathbf{x}_{1},\mathcal{Z},X,Y),\ldots,\pi(\mathbf{x}_{r},\mathcal{Z},X,Y)\}$
7:     $\theta_{G}\leftarrow\theta_{G}||\text{rand()},\theta_{P}\leftarrow\theta_{P}||\text{rand()}$()
8:     Train net $P$ with $\mathcal{R}$ and $\mathcal{R}_{\text{fitness}}$
9:     for each epoch do
10:         $\mathbf{z}_{1},\mathbf{z}_{2}\sim\mathcal{N}(0,1)^{Q}$
11:         $\mathbf{x}_{1}=M(G(\mathbf{z}_{1})),\ \mathbf{x}_{2}=M(G(\mathbf{z}_{2}))$
12:         $\mathcal{L}(\theta_{G})=\mathcal{L}_{G}(\mathbf{z}_{1},\mathbf{z}_{2},\mathbf{x}_{1},\mathbf{x}_{2},\theta_{P})$
13:         $\theta_{G}\leftarrow\text{GradientDescent}(\mathcal{L}(\theta_{G}))$
14:         $\mathcal{Z}_{\text{new}}=\text{parse}(\mathbf{x}_{1})\cup\text{parse}(\mathbf{x}_{2})$
15:         for $f_{\text{new}}$ in $\mathcal{Z}_{\text{new}}$ do
16:              if $f_{\text{new}}$ is qualified and $f_{\text{new}}\notin\mathcal{Z}$ then
17:                  $\mathcal{Z}\leftarrow\mathcal{Z}\cup\{f_{\text{new}}\}$
18:              end if
19:         end for
20:         $\mathcal{R}\leftarrow\mathcal{R}\cup\{\mathbf{x}_{1},\mathbf{x}_{2}\}$
21:     end for
22:end while
23:return $\mathcal{Z}$
24:For demonstration convenience, the batch size dimension is not shown.
3.2. Alpha Combination
@@
Given the potential periodic or permanent ineffectiveness of certain factors due to congestion, shifts in market style, etc., the utilization of fixed factor weights proves inadequate in promptly adjusting to changes in factor weights. This inadequacy can result in overfitting of the training set. The factor metrics serve as performance indicators over a specific time frame. Upon the arrival of new data, the evaluation indicators for factors, including Information Coefficient (IC), IC Information Ratio (ICIR), and RankIC, undergo alterations. Due to the momentum effect observed in factor performance, factors that have demonstrated success in the past tend to exhibit positive performance in the future(EHSANI and LINNAINMAA, 2022).
To address these challenges, we have developed a dynamic weight factor combination model. At each time point $t$, leveraging data from the preceding $n$ days, we conduct a reassessment of the factors within the factor zoo $\mathcal{Z}$. The factors are re-ranked and selected based on their most recent performance metrics, including ICIR, IC, RankIC, etc. Subsequently, we employ the latest data to fit the best linear model for predicting the current combination of ‘N’ factors. This model is then utilized to predict the current data point. The detailed workings of our combinational model are demonstrated in Algorithm 2 and (B) at Figure 2
@@
Algorithm 2 Factor Combining Pipeline
1:Input:Factor zoo $\mathcal{Z}=\{f_{1},\ldots,f_{k}\}$, max factor number $N$, dataset $X=\{X_{t}\},Y=\{y_{t}\}$
2:Output: Prediction $\hat{Y}=\{\hat{y_{t}}\}$
3:$\hat{Y}\xleftarrow{}{\phi}$
4:for $t\leftarrow 1$ to $T$ do
5:     $\mathcal{Z}_{t}={\phi}$
6:     for all $f\in\mathcal{Z}$ do
7:         Calculate $IC_{t}\rho(f),ICIR_{t}\hat{\rho}(f)$
[…]
9:              $\mathcal{Z}_{t}\xleftarrow[]{}\mathcal{Z}_{t}\cup\{f\}$
10:         end if
11:     end for
12:     Sort $\mathcal{Z}_{t}$ based on $IC_{t}\rho(f)$
13:     Select the top $N$ factors from $\mathcal{Z}_{t}$: $\mathcal{Z}_{t}^{(N)}=\text{Top-N}(\mathcal{Z}_{t})$
14:     $Model=LinearRegression(\mathcal{Z}_{t}^{(N)},y_{t})$
15:     $\hat{y_{t}}\xleftarrow[]{}Model.Predict(X_{t})$
16:     $\hat{Y}\xleftarrow[]{}\hat{Y}\cup\hat{y_{t}}$
17:end for
18:return $\hat{Y}$
3.3. Overview
Figure 2. The illustration of our overall framework. (A) Alpha Factor Generating Model which generates the factor zoo with the training set data. (B) Demonstrates the process of combining Mega-Alpha on day t, a process iteratively executed for each trading day.
@@ 3.3. Overview
Once the Alpha Factor Zoo has been extracted,it serves as a fixed input to the combination model and remains unchanged thereafter. During the inference and trading phases, the combination model utilizes updated historical data at each time step $t$ to reassess the recent performance of each factor within the Factor Zoo. Based on this evaluation, the model filters and integrates factors to formulate the Mega-Alpha signal for the given day. The hyperparameters is shown in Appendix B and our framework implementation is published on GitHub [^1] https://github.com/DulyHao/AlphaForge.
4. Experiments
@@
4.1. Experiments settings
4.1.1. Data
The market styles are diverse and ever-changing, potentially leading to over-fitting of models during the training, validation, and testing dataset splits. In practical investment contexts, over-fitting often yields adverse outcomes. Moreover, in real-world investment practices, it is essential to periodically re-calibrate models with the influx of new real-time data. To mitigate over-fitting and closely emulate the actual investment process, we conducted performance testing from 2018 to 2022. The model was retrained annually with updated data, using the year preceding the test year as the validation dataset, resulting in a total of five training sessions. The first training set, validation set and test set are respectively (2010-01-01 to 2016-12-31), (2017-01-01 to 2017-12-31) and (2018-01-01 to 2018-12-31). The last data set segmentation session is (2010-01-01 to 2020-12-31), (2021-01-01 to 2021-12-31) and (2022-01-01 to 2022-12-31).
In practical investment scenarios, the generation of trading signals by the model is contingent upon the closure of the market and the acquisition of closing prices. Predicting returns based on closing prices may not closely align with real-world scenario. Typically, funds with substantial trading volumes execute trades gradually throughout the following day upon receiving the trading signal. The transaction price is often benchmarked against the volume-weighted average price (VWAP) of the next day. Consequently, we define the optimization goal for factor mining as Ref(VWAP, -21)/Ref(VWAP, -1) - 1, representing the return rate of buying the day after receiving the signal and selling 20 days later. While this approach better mirrors real-world scenarios, it may result in a reduction in the value of evaluation indicators. The experimental framework involves utilizing the constituent stocks of CSI300 and CSI500 as the stock pools.
4.1.2. Compared Methods
To assess the distinction of our framework in comparison to traditional formulaic Alpha factor generation methods, we designed two methodologies for comparison against our approach. These include the Genetic Programming (GP) method and the current state-of-the-art method, Reinforcement Learning (RL)(Yu et al., 2023). GP employs the Information Coefficient (IC) as the optimization objective, generating formula trees through genetic approaches. RL utilizes reinforcement learning techniques to generate a set of alpha factors, with the optimization objective being the IC of a Mega-Alpha composed of these factors. To avoid the effects of random seeds, we repeated the run 5 times for each model.
4.1.3. Additional Experiments

### v3

@@ Abstract
The complexity of financial data, characterized by its variability and low signal-to-noise ratio, necessitates advanced methods in quantitative investment that prioritize both performance and interpretability.Transitioning from early manual extraction to genetic programming, the most advanced approach in the alpha factor mining domain currently employs reinforcement learning to mine a set of combination factors with fixed weights. However, the performance of resultant alpha factors exhibits inconsistency, and the inflexibility of fixed factor weights proves insufficient in adapting to the dynamic nature of financial markets. To address this issue, this paper proposes a two-stage formulaic alpha generating framework AlphaForge, for alpha factor mining and factor combination. This framework employs a generative-predictive neural network to generate factors, leveraging the robust spatial exploration capabilities inherent in deep learning while concurrently preserving diversity. The combination model within the framework incorporates the temporal performance of factors for selection and dynamically adjusts the weights assigned to each component alpha factor. Experiments conducted on real-world datasets demonstrate that our proposed model outperforms contemporary benchmarks in formulaic alpha factor mining. Furthermore, our model exhibits a notable enhancement in portfolio returns within the realm of quantitative investment and real money investment.
Introduction
@@
A central challenge in the field of quantitative investment is stock trend forecasting. This difficulty arises from the inherent characteristics of stock data, featuring a low signal-to-noise ratio and considerable noise (Qian, Hua, and Sorensen 2007). Professionals and researchers commonly employ a strategy involving the extraction of Alpha factors from raw historical data to predict future returns. The contemporary approaches to Alpha factor mining can be broadly categorized into two main methodologies: machine learning methods and formulaic Alpha methods. Deep learning models such as LSTM(Hochreiter and Schmidhuber 1997) and Historical Inference Sequence Transformers (HIST)(Xu et al. 2021a), among others, are commonly employed to generate more complex Alphas, which often lack interpret-ability. On the other hand, formulaic Alpha mining aims to identify simple formulas that can replace Alpha factors. Early approaches centered on the manual extraction of factors with economic significance, such as the Fama three-factor model(Fama and French 1992). However, as these factors became widely known and used, their ability to predict outcomes started to decrease(Kakushadze 2016a). Over time, to overcome the limitations of traditional factors, researchers began exploring alternative methods, such as genetic programming, to automatically generate more effective factors. (Kakushadze 2016a) conducts an analysis of 101 formulaic Alpha factors within the US market. Employing a genetic programming algorithm, the study systematically explores individual, unrelated factors through genetic variation of formula trees. The current method involves using reinforcement learning algorithms to simultaneously identify a combination of Alpha factors along with their associated weights(Yu et al. 2023), with the objective of optimizing the discovery of the most robust composite factors.
@@
In investment practice, investment managers often collect a large batch of Alpha factors into a factor library. These alpha factors are combined through a combination model to form ”Mega-Alpha”, a final signal used for trading. With each new data arrival, factor values are recalculated, and Mega-Alpha is computed for trading decisions. Due to the financial market’s emphasis on interpretability, the models for combining factors typically adopt linear structures. Cutting-edge Reinforcement Learning (RL) methods(Yu et al. 2023) integrate the combination model and mining process into a unified framework, determining both factor discovery and their combination into Mega-Alpha. However, despite extensive factor exploration, only a fixed subset is typically utilized in actual investment. Additionally, the cyclic nature of each Alpha factor’s stock selection capability, and the potential for reversal over time, must be considered. Fixed factors and weights sometimes render parts of the Mega-Alpha ineffective or even reverse their effects.
@@
The factor mining model, inspired by (Linder et al. 2020), utilizes a surrogate model to learn the distribution of alpha factor scores. The generative model is trained to maximize the output of the surrogate model, facilitating the generation of high-scoring factors. Gradient-based methods ensure the generation of desirable factors even in the extremely sparse score space, allowing adjustments to the scoring function based on previously mined factors to ensure low correlation.
@@
  3. 3.
    We conducted a comprehensive set of experiments to validate the effectiveness of our proposed methodology. Subsequent additional experiments and real investment provided evidence that factor timing can result in profitable outcomes.
Preliminary
Alpha Factor Definition
@@
Alpha Factor Metrics
@@
$IC(f,X,Y)=\mathbb{E}_{t}\left[\rho\left(v_{t},y_{t}\right)\right]=\frac{1}{T}\sum_{t=1}^{T}\rho\left(v_{t},y_{t}\right)$ (1)
where $v_{t},y_{t}\in\mathbb{R}^{n}$. The correlation $\rho\left(v_{t},y_{t}\right)$ for each cross-section depicts the relationship between the factor value and the subsequent period’s return. The IC describes the overall stock-picking ability of the factor, with higher values indicating superior stock-picking performance. In addition, the inclusion of RankIC is necessary to complement the measurement indicators because of the instability of pearson correlation. More details about these metrics could be found in the supplement materials.
Formulaic Alpha
@@
The representation of formulas traditionally involves expression trees, where operands are leaf nodes, and operators are non-leaf nodes. Initial attempts at formula generation employed genetic programming algorithms on these trees. In alignment with contemporary deep learning methods, our approach follows the methodology introduced in (Yu et al. 2023). This involves representing formulas using Reverse Polish Notation (RPN), acquired through post-order traversal of the formula tree. This facilitates the representation of formulas in a one-hot matrix format compatible with deep learning networks.
Methodology
@@
Factor Mining Model
Algorithm 1 Factor Mining Pipeline
Input: stock data including data and target $X=\{X_{t}\},Y=\{y_{t}\}$
Output: A group of respectively low correlation strong factors which is called factor zoo: $\mathcal{Z}=\{f_{1},\ldots,f_{k}\}$
Initialize the factor zoo $\mathcal{Z}=\emptyset$
Sample a group of randomized factors onehot matrices =$\mathcal{R}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{r}\}$
while $|\mathcal{Z}|<\text{TargetFactorNum}$ do
   $\mathcal{R}_{\text{fitness}}=\{\pi(\mathbf{x}_{1},\mathcal{Z},X,Y),\ldots,\pi(\mathbf{x}_{r},\mathcal{Z},X,Y)\}$
   $\theta_{G}\leftarrow\theta_{G}||\text{rand()},\theta_{P}\leftarrow\theta_{P}||\text{rand()}$()
   Train net $P$ with $\mathcal{R}$ and $\mathcal{R}_{\text{fitness}}$
   for each epoch do
      $\mathbf{z}_{1},\mathbf{z}_{2}\sim\mathcal{N}(0,1)^{Q}$
      $\mathbf{x}_{1}=M(G(\mathbf{z}_{1})),\ \mathbf{x}_{2}=M(G(\mathbf{z}_{2}))$
      $\mathcal{L}(\theta_{G})=\mathcal{L}_{G}(\mathbf{z}_{1},\mathbf{z}_{2},\mathbf{x}_{1},\mathbf{x}_{2},\theta_{P})$
      $\theta_{G}\leftarrow\text{GradientDescent}(\mathcal{L}(\theta_{G}))$
      $\mathcal{Z}_{\text{new}}=\text{parse}(\mathbf{x}_{1})\cup\text{parse}(\mathbf{x}_{2})$
      for $f_{\text{new}}$ in $\mathcal{Z}_{\text{new}}$ do
         if $f_{\text{new}}$ is qualified and $f_{\text{new}}\notin\mathcal{Z}$ then
            $\mathcal{Z}\leftarrow\mathcal{Z}\cup\{f_{\text{new}}\}$
         end if
      end for
      $\mathcal{R}\leftarrow\mathcal{R}\cup\{\mathbf{x}_{1},\mathbf{x}_{2}\}$
   end for
end while
return $\mathcal{Z}$
For demonstration convenience, the batch size dimension is not shown.
@@
$\mathcal{L}_{P}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left(P\left(\boldsymbol{x}_{i}\right)-\text{ fitness }\left(\boldsymbol{x}_{\boldsymbol{i}}\right)\right)^{2}}.$ (2)
@@
$\mathcal{L}_{\text{Fitness }}=-P(M(G(z)))$ (3)
@@
However, focusing solely on optimizing for high fitness may lead to premature convergence of the network $G$ to a local optimum. Therefore, it is necessary to introduce a diversity loss to force the generator $G$ to produce a diverse array of alpha factor formulas. To achieve this, we introduce a loss by generating two sets of factors based on two samplings of $z_{1}$ and $z_{2}$, subsequently penalizing the correlation between these two sets of factors. The final loss function for training the generator $\mathrm{G}$ is:
{dmath}
L_G=L_Fitness +L_Diversity =-P(x_1)+λ_onehot * Similarity _onehot (f(z_1), f(z_2))+λ_hidden * Similarity _hidden (f(z_1), f(z_2))
@@
Our framework is designed to identify a set of highly effective strong factors in the factor mining stage for the factor timing combination model. This involves an investigation, including a comparison between strong and weak factor pools. A strict criterion is maintained to ensure that factors included in the library meet specific requirements. Incorporating domain knowledge, our criteria for factor entry comprise three fundamental aspects:IC and ICIR are used to filter the stock-picking ability and stability, the correlation of returns with existing factors in the library avoids overlapping with the stock-picking ability of existing factors. Due to the implementation of a generative-predictive architecture coupled with gradient-based algorithms, the generator is capable of capturing the essential characteristics of factors in a ”directional” manner, even under conditions of significant sparsity within the fitness function: {dmath} π(x, Z, X, Y)= { Abs (IC(f, X, Y)),f is valid and ψ(f, Z, X, Y)¡CORR0, else Where $f=\operatorname{parse}(x)$ represents the operational formula parsed from the onehot matrix representation $x$, and $\mathcal{Z}$ denotes the existing factor set which is called factor zoo. The absolute value of IC is taken because a negative IC factor can be transformed into a positive IC factor by reversing the value of the factor. The $\psi$ function computes the maximum absolute value of the correlation between $f$ and each existing factor in $\mathcal{Z}$. Here, CORR’ is a manually set parameter. When $|Z|=0$, $\pi(f,\mathcal{Z},X,Y)$ returns the absolute value of the factor’s IC. Algorithm 1 provides a detailed description of our alpha mining model.
Figure 1: The illustration of our overall framework. (A) Alpha Factor Generating Model which generates the factor zoo. (B) Demonstrates the process of combining Mega-Alpha on day t, a process iteratively executed for each trading day.
Alpha Combination
Algorithm 2 Factor Combining Pipeline
Input:Factor zoo $\mathcal{Z}=\{f_{1},\ldots,f_{k}\}$, max factor number $N$, dataset $X=\{X_{t}\},Y=\{y_{t}\}$
Output: Prediction $\hat{Y}=\{\hat{y_{t}}\}$
$\hat{Y}\xleftarrow{}{\phi}$
for $t\leftarrow 1$ to $T$ do
   $\mathcal{Z}_{t}={\phi}$
   for all $f\in\mathcal{Z}$ do
      Calculate $IC_{t}\rho(f),ICIR_{t}\hat{\rho}(f)$
[…]
         $\mathcal{Z}_{t}\xleftarrow[]{}\mathcal{Z}_{t}\cup\{f\}$
      end if
   end for
   Sort $\mathcal{Z}_{t}$ based on $IC_{t}\rho(f)$
   Select the top $N$ factors from $\mathcal{Z}_{t}$: $\mathcal{Z}_{t}^{(N)}=\text{Top-N}(\mathcal{Z}_{t})$
   $Model=LinearRegression(\mathcal{Z}_{t}^{(N)},y_{t})$
   $\hat{y_{t}}\xleftarrow[]{}Model.Predict(X_{t})$
   $\hat{Y}\xleftarrow[]{}\hat{Y}\cup\hat{y_{t}}$
end for
return $\hat{Y}$
@@
Given the potential periodic or permanent ineffectiveness of certain factors due to congestion, shifts in market style, etc., the utilization of fixed factor weights proves inadequate in promptly adjusting to changes in factor weights. This inadequacy can result in overfitting of the training set. The factor metrics serve as performance indicators over a specific time frame. Upon the arrival of new data, the evaluation indicators for factors undergo alterations. Due to the momentum effect observed in factor performance, factors that have demonstrated success in the past tend to exhibit positive performance in the future(EHSANI and LINNAINMAA 2022).
To address these challenges, we have developed a dynamic weight factor combination model. At each time point $t$, leveraging data from the preceding $n$ days, we conduct a reassessment of the factors within the factor zoo $\mathcal{Z}$. The factors are re-ranked and selected based on their most recent performance metrics, including ICIR, IC, RankIC, etc. Subsequently, we employ the latest data to fit the best linear model for predicting the current combination of ‘N’ factors. This model is then utilized to predict the current data point. The detailed workings of our combinational model are demonstrated in Algorithm 2 and (B) at Figure 1 .
@@
Overview
@@
Once the Alpha Factor Zoo has been extracted,it serves as a fixed input to the combination model and remains unchanged thereafter. During the inference and trading phases, the combination model utilizes updated historical data at each time step $t$ to reassess the recent performance of each factor within the Factor Zoo. Based on this evaluation, the model filters and integrates factors to formulate the Mega-Alpha signal for the given day. The hyperparameters could be found in supplementary materials and our framework implementation is published on GitHub [^1] https://github.com/DulyHao/AlphaForge.
Experiments
@@
Experiments settings
Data
We choose the CSI300 and CSI500 dataset because the constituent stocks of these two indices cover the majority of the market capitalization in China’s A-share market. Additionally, the studies we referenced also focused on these two datasets(Yu et al. 2023; Xu et al. 2021a; Yang et al. 2020). The market styles are diverse and ever-changing, potentially leading to over-fitting of models during the training, validation, and testing dataset splits. In practical investment contexts, over-fitting often yields adverse outcomes. Moreover, in real-world investment practices, it is essential to periodically re-calibrate models with the influx of new real-time data. To mitigate over-fitting and closely emulate the actual investment process, we conducted performance testing from 2018 to 2022. The model was retrained annually with updated data, using the year preceding the test year as the validation dataset, resulting in a total of five training sessions. The first training set, validation set and test set are respectively (2010-01-01 to 2016-12-31), (2017-01-01 to 2017-12-31) and (2018-01-01 to 2018-12-31). We use ’Ref(VWAP, -21)/Ref(VWAP, -1) - 1’ as the label because it more closely reflects real-world scenarios, although it may lead to changes in the metrics. More details on this can be found in the supplementary materials. The stock data is public data and we use Qlib(Yang et al. 2020) to download it.
Compared Methods
To assess the distinction of our framework in comparison to traditional formulaic Alpha factor generation methods, we designed three methodologies for comparison against our approach. These include the Genetic Programming (GP) method, the Deep Symbolic Optimization DSO method (Landajuela et al. 2022a) and Reinforcement Learning (RL)(Yu et al. 2023). GP employs the Information Coefficient (IC) as the optimization objective, generating formula trees through genetic approaches. As a representative method of symbolic regression, the DSO method aligns best with our task. RL utilizes reinforcement learning techniques to generate a set of alpha factors, with the optimization objective being the IC of a Mega-Alpha composed of these factors. To avoid the effects of random seeds, we repeated the run 5 times for each model.
We further incorporated three Machine Learning based models for benchmarking purposes: XGBoost(Chen and Guestrin 2016), an ensemble learning method using gradient boosting with decision trees, effective for capturing non-linear relationships in stock market data; LightGBM(Ke et al. 2017), a highly efficient gradient boosting framework, optimized for handling large-scale financial datasets; and MLP (Multilayer Perceptron), a neural network model capable of learning complex patterns in financial data through multiple interconnected layers of nodes.
Additional Experiments